OpenStudy (anonymous):

write an equation of the circle with the given center and radius center: (-7,-6) and radius 2

hartnn (hartnn):

Equation of circle with center (h,k) and radius 'r' is \(\Large (x-h)^2+(y-k)^2 = r^2\)

hartnn (hartnn):

here, h,k = -7,-6 r=2 just plug in!

OpenStudy (anonymous):

Okay so it'll be like this: (x-(-7)^+(y-(6)^=2^

OpenStudy (anonymous):

Is this right?

hartnn (hartnn):

yes whats \(\Large x-(-7) = ...?\)

OpenStudy (anonymous):

'x' minus -7.

hartnn (hartnn):

two negatives become a positive!

OpenStudy (anonymous):

with the two negatives means the '7 turns into positive 7 right?

hartnn (hartnn):

yess \(x-(-7) = x+7\) \(y- (-6) = ...?\)

OpenStudy (anonymous):

y+6 right? So it would be (x+7)^ + (y+6)^=2 right?

hartnn (hartnn):

you forgot to square the radius ? \((x+7)^2 + (y+6)^2 = 2^2 \\(x+7)^2 + (y+6)^2 = 4 \)

OpenStudy (anonymous):

Okay so now I don't know what to do! What should I do next>

hartnn (hartnn):

thats the equation of your circle! thats it!

OpenStudy (anonymous):

OOOOOOOOOH! Okay then! Thanks so much for the help again! You're a real life saver! I totally appreciate it!

hartnn (hartnn):

you're most welcome ^_^